# Document-Level In-Context Few-Shot Relation Extraction via Pre-Trained Language Models

### v2

@@ Abstract
Relation extraction aims at inferring structured human knowledge from textual documents. State-of-the-art methods based on language models commonly have two limitations: (1) they require named entities to be either given as input or infer them, which introduces additional noise, and (2) they require human annotations of documents. As a remedy, we present a novel framework for document-level in-context few-shot relation extraction via pre-trained language models. We achieve crucial benefits in that we eliminate the need for both named entity recognition and human annotation of documents. Unlike existing methods based on fine-tuning, our framework is flexible in that it can be easily updated for a new set of relations without re-training. We evaluate our framework using DocRED, the largest publicly available dataset for document-level relation extraction, and demonstrate that our framework achieves state-of-the-art performance. Finally, we show that our framework actually performs much better than the original labels from the development set of DocRED. To the best of our knowledge, we are the first to reformulate the document-level relation extraction task as a tailored in-context few-shot learning paradigm.
LLM, relation extraction, in-context learning
[…]
*supplements
@@ 1 Introduction
Relational facts are widely used to represent human knowledge (Grishman, 2019; Han et al., 2020; Weikum et al., 2021). With the explosion of the web, relational facts have become broadly available through large knowledge bases (KBs) (Auer et al., 2007; Suchanek et al., 2007; Bollacker et al., 2008; Vrandečić and Krötzsch, 2014) and thereby support many downstream tasks. Examples are commonsense reasoning (Lin et al., 2019; Liu et al., 2021a), question answering (Luo et al., 2018; Das et al., 2022; Wang et al., 2022), fact checking (Huynh and Papotti, 2019; Vedula and Parthasarathy, 2021), and product recommendations (Wang et al., 2018; Zhou et al., 2020a, b). However, relational facts are not readily available in structured form but are commonly embedded in unstructured texts. To this end, methods are needed for relation extraction from text.
State-of-the-art methods for relation extraction leverage pre-trained language models (LMs) and fine-tune them using human-annotated documents. These works can be loosely grouped into two streams. One stream requires named entities (e. g., by specifying the entities of interest with a special token) as input (Wang et al., 2019; Xu et al., 2021; Zhang et al., 2021; Zhou et al., 2021; Tan et al., 2022; Wang Xu and Zhao, 2022; Xiao et al., 2022; Hu et al., 2023; Xu et al., 2023). Another stream avoids the use of named entities as input and, instead, learns the detection of named entities through a tailored training objective (Cabot and Navigli, 2021; Eberts and Ulges, 2021; Giorgi et al., 2022; Lu et al., 2022b; Zhang et al., 2023b).
Yet, state-of-the-art methods based on LMs have three main drawbacks that limit their applicability for relation extraction in practice. (1) State-of-the-art methods for relation extraction typically require the named entities to be either given as input or to be inferred via a customized training objective. This introduces additional noise in the relation extraction pipeline and thus hinders the downstream performance. (2) State-of-the-art methods for relation extraction need large amounts of human-annotated documents for training. However, human annotation is costly. (3) State-of-the-art methods are based on LMs that are fine-tuned. As a result, whenever a new type of relation is added to the knowledge base or whenever a better LM is adopted, the entire training process must be repeated. This introduces a huge computational overhead.
There are some recent efforts that use the reasoning abilities of LMs via in-context learning for relation extraction (Wadhwa et al., 2023; Wan et al., 2023). However, these are designed for sentence-level relation extraction, meaning for a small set of relation types. Hence, due to high computational costs, their scalability to documents is limited (see Table 1). Here, we introduce a novel method to leverage in-context learning for document-level relation extraction.
@@
Contributions: [^1] Codes are available at https://github.com/oezyurty/REPLM .
  1. 1.
    We present a novel framework called REPLM for in-context few-shot relation extraction via pre-trained LMs. To the best of our knowledge, we are the first to reformulate the document-level relation extraction task as a tailored in-context few-shot learning paradigm.
  2. 2.
    Our REPLM framework has key advantages for practice: it eliminates the noise from named entity recognition, it circumvents the need for human annotations, and it is flexible in that it is directly applicable to new relations without re-training.
  3. 3.
    We show that our REPLM framework achieves state-of-the-art performance.
@@ 2 Related Work
In-context few-shot learning of LMs: LMs have achieved superior performance in many downstream tasks (Beltagy et al., 2019; Devlin et al., 2019; Liu et al., 2019; Radford et al., 2019; Brown et al., 2020; Lewis et al., 2020; Raffel et al., 2020; Wang and Komatsuzaki, 2021; Wei et al., 2022a, b; OpenAI, 2023; Wang et al., 2023; Zhang et al., 2023c). Due to the large computational cost of fine-tuning an LM, Brown et al. (2020) proposed in-context few-shot learning paradigm to teach an LM a new task at inference time. We provide an overview of applications in Appendix A. However, we are not aware of any earlier work that leveraged the in-context few-shot learning paradigm for document-level relation extraction.
@@
Early research on relation extraction: Early works extracted relations from text via pattern extraction methods (Carlson et al., 2010; Nakashole et al., 2012; Jiang et al., 2017; Pawar et al., 2017; Weikum et al., 2021) and via statistical methods (Sarawagi and Cohen, 2004; Zhang et al., 2006a, b; Jiang and Zhai, 2007; Nguyen et al., 2007; Wang, 2008; Yu and Lam, 2010; Wang et al., 2014; Lin et al., 2015). However, the above methods have only a limited modeling capacity, as compared to neural networks (Zeng et al., 2014; Miwa and Bansal, 2016; Zhou et al., 2016; Adel and Schütze, 2017; Katiyar and Cardie, 2017; Zheng et al., 2017; Han et al., 2020). As shown later, LM-based methods better capture the complex interactions between named entities to classify the relation. A detailed review is in the Appendix A.
@@
Relation extraction via LM: State-of-the-art methods for relation extraction are based on fine-tuning pre-trained LMs. Specifically, these methods use pre-trained LMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and SciBERT (Beltagy et al., 2019) and fine-tuned them for relation extraction. For instance, Wang et al. (2019) fine-tuned BERT to classify the relation between each named entity pair in a sentence. There have been various follow-up works to improve performance by learning complex dependency between named entities (Wang and Lu, 2020; Paolini et al., 2021; Xu et al., 2021; Zhang et al., 2021; Zhou et al., 2021; Tan et al., 2022; Wang Xu and Zhao, 2022; Xiao et al., 2022; Hu et al., 2023; Xu et al., 2023; Zhang et al., 2023a). A detailed review is in Appendix A. However, all these works require the named entities to be annotated and provided as input at both training and test time.
@@
Some works relax the requirement of given named entities to facilitate processing the raw documents at test time. As a remedy, these works jointly learn to extract named entities and relations. Examples are SpERT (Eberts and Ulges, 2020), JEREX (Eberts and Ulges, 2021), Seq2Rel (Giorgi et al., 2022), UIE (Lu et al., 2022b), and TaG (Zhang et al., 2023b). Their drawback is that multi-step pipelines with named entities recognition propagate the errors to relation extraction (Cabot and Navigli, 2021). Motivated by this, Cabot and Navigli (2021) developed REBEL, an auto-regressive model based on BART (Lewis et al., 2020), which is fine-tuned to output relations as sequences of texts. To the best of our knowledge, REBEL is the only LM-based method without the need for a named entity recognition pipeline and thus represents our main baseline.
@@
Wan et al. (2023) Sentence OpenAI’s GPT
@@
REBEL (Cabot et al., 2021) Document BART-large
@@
In-context learning for sentence-level relation extraction:. There are two recent works that leverage in-context learning for sentencel-level relation extraction (see Table 1).Wan et al. (2023) requires named entities to be provided for each sentence and generates $k$ different chain-of-thought (CoT) reasonings (Wei et al., 2022b) from GPT-3 (Brown et al., 2020) as in-context examples for each named entity pairs to classify the relation between them. Similarly, Wadhwa et al. (2023) requires CoT reasonings from GPT-3 for the entire training corpus and fine-tunes Flan-T5 (Chung et al., 2022) based on the generated CoT outputs. However, these works are only applicable to a small set of sentences and relation types due to their high computational cost resulting from commercial architectures. Because of that, these methods work only at the sentence-level relation extraction (i. e., they are not scalable to document-level).
@@ 3 Problem Description
Relation extraction: The task of relation extraction from the documents is typically defined as follows (Wang et al., 2019; Eberts and Ulges, 2021; Xu et al., 2021; Zhang et al., 2021; Zhou et al., 2021; Giorgi et al., 2022; Lu et al., 2022b; Tan et al., 2022; Wang Xu and Zhao, 2022; Xu et al., 2023). Given is a set of documents $\mathcal{D}=\{d_{i}\}_{i=1}^{M}$, where $M$ is the number of documents in the set. For each document $d_{i}$, the aim is to enumerate knowledge triplets $\{(r_{im},s_{im},o_{im})\}_{m=1}^{R_{i}}$, where $r_{im}\in\mathcal{R}$ is a relation and $s_{im}$ and $o_{im}$ are the subject and object of the relation $r_{im}$, and where $R_{i}$ is the number of relations in $d_{i}$. For instance, the document “The Reality Dysfunction is a science fiction novel by British writer Peter F. Hamilton normal-$\ldots$” yields the knowledge triplets (author, The Reality Dysfunction, Peter F. Hamilton), (genre, The Reality Dysfunction, science fiction), etc.
Difference to earlier works: Earlier works (see Sec. 2) generally address the above task through an intermediate step that involves named entities detection. Specifically, the aforementioned works first detect the named entities of a document $d_{i}$, i. e., $\{e_{ij}\}_{j=1}^{N_{i}}$, where $N_{i}$ is the number of entities in $d_{i}$. Then they proceed by predicting the relation(s) between each named entity pair $(e_{ij},e_{ij^{\prime}})_{j,j^{\prime}\in\{1,\ldots,N_{i}\},j\neq j^{\prime}}$ among the $R$ relations, where $e_{ij}$ is the subject and $e_{ij^{\prime}}$ is the object of predicted relation(s). As such, the number of predictions scales with the number of named entity pairs, i.e., it is in $O(N_{i}^{2})$. For the prediction, state-of-the-art methods make use of pre-trained LMs for which a multi-class classification head is fine-tuned. The classification head also predicts when there is no relation between the given pair and then discards the knowledge triplet. All remaining relations give the knowledge triplets $\{(r_{im},s_{im},o_{im})\}_{m=1}^{R_{i}}$.
@@ 4 Proposed REPLM Framework
Approach (see Fig. 1): At a high-level, our framework seeks to infer the correct knowledge triplets $(r,s,o)$ from a given document $d_{i}$ and for a given relation $r$. To do so, we estimate the joint probability of a subject-object pair $(s,o)$ conditional on $d_{i}$ and $r$, i.e., $p(s,o\,|\,d_{i},r)$. After having estimated the probability, we simply rank the candidate subject-object pairs according to their probabilities and keep the top-ranked pairs as knowledge triplets. In our framework, we follow this approach but, as a main innovation, leverage a pre-trained LM to approximate $p(s,o\,|\,d_{i},r)$.
However, selecting only a single set of in-context examples may lead to a poor approximation of the probability $p(s,o\,|\,d_{i},r)$, because the selected in-context examples introduce bias in output generation (e. g., recency bias, label space of the in-context examples) as studied in prior literature (Liu et al., 2022a; Min et al., 2022b; Rubin et al., 2022; Hongjin et al., 2023; Wei et al., 2023). Instead, we mitigate the above bias by considering multiple sets of in-context examples. As a result, we calculate the joint probability of a subject-object pair as
@@ 4.1 Selecting Candidates for In-Context Examples
Distant supervision assumes that, if a document $\tilde{d}\in\mathcal{D}^{\text{dist}}$ contains both the subject and object of a knowledge triplet from a KB, it likely discusses their relationship. This premise allows for the automatic generation of annotated document sets. Such automated annotation introduces noisy labels (Yao et al., 2019; Cabot and Navigli, 2021), yet that can oftentimes be effective (Min et al., 2022b). A key benefit is that distantly supervised documents offer rich insights into label space, textual distributions, and expected output formats, over which the in-context few-shot learning paradigm in our REPLM can generalize. [^2] We further compare distant supervision with human-annotated data. They have the same performance, confirming the effectiveness of this approach for relation extraction (see Appendix D).
Semantic filtering: (i) We first filter the documents in $\mathcal{D}^{\text{dist}}$ so that we only keep the documents that contain at least one knowledge triplet of a relation $r$. We denote the result by $\mathcal{D}_{r}^{\text{dist}}$. Formally, $\mathcal{D}_{r}^{\text{dist}}$ is defined as
$\mathcal{D}_{r}^{\text{dist}}=\{d_{j}\,|\,\exists\,r^{\prime},s,o\ \mathrm{s.t.}\ (r^{\prime},s,o)\in d_{j}\\ \land r^{\prime}=r\land d_{j}\in\mathcal{D}^{\text{dist}}\}.$

### v4

@@ Abstract
Document-level relation extraction aims at inferring structured human knowledge from textual documents. State-of-the-art methods for this task use pre-trained language models (LMs) via fine-tuning, yet fine-tuning is computationally expensive and cannot adapt to new relation types or new LMs. As a remedy, we leverage the generalization capabilities of pre-trained LMs and present a novel framework for document-level in-context few-shot relation extraction. Our framework has three strengths: it eliminates the need (1) for named entity recognition and (2) for human annotations of documents, and (3) it can be updated to new LMs without re-training. We evaluate our framework using DocRED, the largest publicly available dataset for document-level relation extraction, and demonstrate that our framework achieves state-of-the-art performance. We further show that our framework actually performs much better than the original labels from the development set of DocRED. Finally, we conduct an extensive benchmark demonstrating the effectiveness of our framework, achieving state-of-the-art results across six relation extraction datasets and outperforming more than 30 baseline methods. Unlike our framework, the baseline methods have large computational overhead (e.g., from fine-tuning). To the best of our knowledge, we are the first to reformulate the document-level relation extraction task as a tailored in-context few-shot learning paradigm.
@@ 1 Introduction
Relational facts are widely used to represent human knowledge (Grishman, 2019, Han et al., 2020, Weikum et al., 2021). With the explosion of the web, relational facts have become broadly available through large knowledge bases (KBs) (Auer et al., 2007, Bollacker et al., 2008, Suchanek et al., 2007, Vrandečić & Krötzsch, 2014) and thereby support many downstream tasks. Examples are commonsense reasoning (Lin et al., 2019, Liu et al., 2021a), question answering (Das et al., 2022, Luo et al., 2018, Wang et al., 2022), fact checking (Huynh & Papotti, 2019, Vedula & Parthasarathy, 2021), and product recommendations (Wang et al., 2018, Zhou et al., 2020a; b). However, relational facts are not readily available in structured form but are commonly embedded in unstructured texts. To this end, methods are needed for relation extraction from text.
State-of-the-art methods for relation extraction leverage pre-trained language models (LMs) and fine-tune them using human-annotated documents. These works can be loosely grouped into two streams. One stream requires named entities (e. g., by specifying the entities of interest with a special token) as input (Hu et al., 2023, Tan et al., 2022, Wang et al., 2019, Wang Xu & Zhao, 2022, Xiao et al., 2022, Xu et al., 2021a; 2023, Zhang et al., 2021, Zhou et al., 2021). Another stream avoids the use of named entities as input and, instead, learns the detection of named entities through tailored training (Cabot & Navigli, 2021, Eberts & Ulges, 2021, Giorgi et al., 2022, Lu et al., 2022b, Zhang et al., 2023b).
Yet, state-of-the-art methods based on LMs have three main drawbacks that limit their applicability for relation extraction in practice. (1) State-of-the-art methods for relation extraction typically require the named entities to be either given as input or to be inferred via a customized training objective. This can propagate the errors into the relation extraction pipeline and thereby degrade the downstream performance. (2) State-of-the-art methods for relation extraction need large amounts of human-annotated documents for training. However, human annotation is costly. (3) State-of-the-art methods are based on LMs that are fine-tuned. As a result, whenever a new type of relation is added to the knowledge base or whenever a better LM is adopted, the entire training process must be repeated. This introduces a huge computational overhead.
There are some recent efforts that use the reasoning abilities of LMs via in-context learning for relation extraction (Li et al., 2023, Wadhwa et al., 2023, Wan et al., 2023). However, these are designed for sentence-level relation extraction, meaning for a small set of relation types. Due to high computational costs, their scalability to documents is limited (see Table 1). Here, we introduce a novel method to leverage in-context learning for document-level relation extraction.
@@
Contributions: [^1] Codes available at https://github.com/oezyurty/REPLM . 1 We present a novel framework called REPLM for in-context few-shot relation extraction via pre-trained LMs. To the best of our knowledge, we are the first to reformulate the document-level relation extraction task as a tailored in-context few-shot learning paradigm. 2 Our REPLM framework has key advantages for practice: it eliminates the error propagation from named entity recognition, it circumvents the need for human annotations, and it is flexible in that it is directly applicable to new relations and new backbone LMs without re-training. 3 We show that our REPLM achieves state-of-the-art performance across a variety of datasets.
@@ 2 Related Work
In-context few-shot learning of LMs: LMs have achieved superior performance in many downstream tasks (Beltagy et al., 2019, Brown et al., 2020, Devlin et al., 2019, Lewis et al., 2020, Liu et al., 2019, OpenAI, 2023, Radford et al., 2019, Raffel et al., 2020, Wang & Komatsuzaki, 2021, Wang et al., 2023b, Wei et al., 2022a; b, Zhang et al., 2023c). Due to the large computational cost of fine-tuning an LM, Brown et al. (2020) proposed in-context few-shot learning to teach an LM a new task at inference time. We provide an overview of applications in Appendix A. However, we are not aware of any earlier work that leveraged in-context few-shot learning for document-level relation extraction.
@@
Early research on relation extraction: Early works extracted relations from text via pattern extraction methods (Carlson et al., 2010, Jiang et al., 2017, Nakashole et al., 2012, Pawar et al., 2017, Weikum et al., 2021) and via statistical methods (Jiang & Zhai, 2007, Lin et al., 2015, Nguyen et al., 2007, Sarawagi & Cohen, 2004, Wang, 2008, Wang et al., 2014, Yu & Lam, 2010, Zhang et al., 2006a; b). However, the above methods have only a limited modeling capacity, as compared to neural networks (Adel & Schütze, 2017, Han et al., 2020, Katiyar & Cardie, 2017, Miwa & Bansal, 2016, Zeng et al., 2014, Zheng et al., 2017, Zhou et al., 2016). As shown later, LM-based methods better capture the complex interactions between named entities to classify the relation. A detailed review is in the Appendix A.
@@
Relation extraction via LM: State-of-the-art methods for relation extraction are based on fine-tuning pre-trained LMs. Specifically, these methods use pre-trained LMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and SciBERT (Beltagy et al., 2019) and fine-tuned them for relation extraction. For instance, Wang et al. (2019) fine-tuned BERT to classify the relation between each named entity pair in a sentence. There have been various follow-up works to improve performance by learning complex dependency between named entities (Hu et al., 2023, Paolini et al., 2021, Tan et al., 2022, Wang & Lu, 2020, Wang Xu & Zhao, 2022, Xiao et al., 2022, Xu et al., 2021a; 2023, Zhang et al., 2023a; 2021, Zhou et al., 2021). A detailed review is in Appendix A. However, these works require the named entities to be annotated and provided as input at both training and test time.
@@
Some works relax the requirement of given named entities to facilitate processing the raw documents at test time. As a remedy, these works jointly learn to extract named entities and relations. Examples are SpERT (Eberts & Ulges, 2020), JEREX (Eberts & Ulges, 2021), Seq2Rel (Giorgi et al., 2022), UIE (Lu et al., 2022b), and TaG (Zhang et al., 2023b). Their drawback is that multi-step pipelines with named entities recognition propagate the errors to relation extraction (Cabot & Navigli, 2021). Motivated by this, Cabot & Navigli (2021) developed REBEL, an auto-regressive model based on BART (Lewis et al., 2020), which is fine-tuned to output relations as sequences of texts. To the best of our knowledge, REBEL is the only fine-tuned LM-based method without the need for a named entity recognition pipeline and thus represents one of our main baselines.
@@
In-context learning for sentence-level relation extraction:. There are three recent works that leverage in-context learning for sentencel-level relation extraction (see Table 1). GPT-RE (Wan et al., 2023) requires named entities to be provided for each sentence and generates $k$ different chain-of-thought (CoT) reasonings (Wei et al., 2022b) from GPT-3 (Brown et al., 2020) as in-context examples for each named entity pairs to classify their relation. CodeIE (Li et al., 2023) leverages Codex (deprecated) model from OpenAI for structured output. It requires that $k$ code generation examples are provided in-context for each relation type. Similar to GPT-RE, Wadhwa et al. (2023) uses CoT reasonings from GPT-3 for the entire training corpus and fine-tunes Flan-T5 (Chung et al., 2022) based on the generated CoT outputs. At the inference time, all of these works require $\mathcal{O}(k\cdot R)$ examples to be fit in-context for $k$-shot demonstrations of $R$ relation types. As a result, these works are only applicable to a small set of sentences and relation types for two reasons: (1) the high computational cost resulting from commercial architectures and (2) the requirement of a large number of in-context examples at the inference time. Therefore, these methods work only for the sentence-level relation extraction (i. e., they are not scalable to document-level).
@@
GPT-RE Wan et al. (2023) Sentence OpenAI’s GPT
CodeIE Li et al. (2023) Sentence OpenAI’s Codex   ✗
@@
REBEL Cabot & Navigli (2021) Document BART-large
@@
Codex models are deprecated at the time of writing.
Our work can easily be extended to other LMs as shown in Section  8.
@@ 3 Problem Description
Relation extraction: The relation extraction from documents is defined as follows (Eberts & Ulges, 2021, Giorgi et al., 2022, Lu et al., 2022b, Tan et al., 2022, Wang et al., 2019, Wang Xu & Zhao, 2022, Xu et al., 2021a; 2023, Zhang et al., 2021, Zhou et al., 2021). Given is a set of documents $\mathcal{D}=\{d_{i}\}_{i=1}^{M}$, where $M$ is the number of documents . For each document $d_{i}$, the aim is to enumerate knowledge triplets $\{(r_{im},s_{im},o_{im})\}_{m=1}^{R_{i}}$, where $r_{im}\in\mathcal{R}$ is a relation and $s_{im}$ and $o_{im}$ are the subject and object of the relation $r_{im}$, and where $R_{i}$ is the number of relations in $d_{i}$. For instance, the document “The Reality Dysfunction is a science fiction novel by British writer Peter F. Hamilton $\ldots$” yields the knowledge triplets (author, The Reality Dysfunction, Peter F. Hamilton), (genre, The Reality Dysfunction, science fiction), etc.
Difference to earlier works: Earlier works (see Sec. 2) generally address the above task through a mandatory step for named entities detection. Specifically, the aforementioned works first need to detect the named entities of a document $d_{i}$, i. e., $\{e_{ij}\}_{j=1}^{N_{i}}$, where $N_{i}$ is the number of entities in $d_{i}$. Then they proceed by predicting the relation(s) between each named entity pair $(e_{ij},e_{ij^{\prime}})_{j,j^{\prime}\in\{1,\ldots,N_{i}\},j\neq j^{\prime}}$ among the $R$ relations, where $e_{ij}$ is the subject and $e_{ij^{\prime}}$ is the object of predicted relation(s). As such, the number of predictions scales with the number of named entity pairs, i.e., it is in $O(N_{i}^{2})$.
@@ 4 Proposed REPLM Framework
Approach (see Fig. 1): At a high level, our framework seeks to infer the correct knowledge triplets $(r,s,o)$ from a given document $d_{i}$ and for a given relation $r$. To do so, we estimate the joint probability of a subject-object pair $(s,o)$ conditional on $d_{i}$ and $r$, i.e., $p(s,o\,|\,d_{i},r)$. After having estimated the probability, we simply rank the candidate subject-object pairs according to their probabilities and keep the top-ranked pairs as knowledge triplets. In our framework, we follow this approach but, as a main innovation, leverage a pre-trained LM to approximate $p(s,o\,|\,d_{i},r)$.
@@
However, selecting only a single set of in-context examples may lead to a poor approximation of the probability $p(s,o\,|\,d_{i},r)$, because the selected in-context examples introduce bias in output generation (e. g., recency bias, label space of the in-context examples) as studied in prior literature (Hongjin et al., 2023, Liu et al., 2022a, Min et al., 2022b, Rubin et al., 2022, Wei et al., 2023). Instead, we mitigate the above bias by considering multiple sets of in-context examples. As a result, we calculate the joint probability of a subject-object pair as
@@ 4.1 Selecting Candidates for In-Context Examples
Distant supervision assumes that, if a document $\tilde{d}\in\mathcal{D}^{\text{dist}}$ contains both the subject and object of a knowledge triplet from a KB, it likely discusses their relationship. This premise allows for the automatic generation of annotated document sets. A key benefit is that distantly supervised documents offer rich insights into label space, textual distributions, and expected output formats, over which the in-context few-shot learning paradigm in our REPLM can generalize. [^2] We further compare distant supervision with human-annotated data. They have the same performance, confirming the effectiveness of this approach for relation extraction (see Appendix D).
Semantic filtering: (i) We first filter the documents in $\mathcal{D}^{\text{dist}}$ so that we only keep the documents that contain at least one knowledge triplet of a relation $r$. We denote the result by $\mathcal{D}_{r}^{\text{dist}}$, defined as
$\mathcal{D}_{r}^{\text{dist}}=\{d_{j}\,|\,\exists\,r^{\prime},s,o\ \mathrm{s.t.}\ (r^{\prime},s,o)\in d_{j}\land r^{\prime}=r\land d_{j}\in\mathcal{D}^{\text{dist}}\}.$ (2)